A 300g mass hangs at the end of a string. A second string hangs from the bottom of that mass and support a 900g mass. (A) find the tension in each string when the masses are accelerating upward at 0.700m/s². (b) Find the tension in each string when the masses are accelerating downward at 0.700m/s².

1
Expert's answer
2021-12-02T10:03:39-0500

(a) Find the tension in each string when the masses are accelerating upward at 0.700 m/s²:

"T_a=(m_a+m_b)(g+a)=12.6\text{ N}.\\
T_b=m_b(g+a)=9.45\text{ N}."

(b) Find the tension in each string when the masses are accelerating downward at 0.700 m/s²:


"T_a=(m_a+m_b)(g-a)=10.9\text{ N}.\\
T_b=m_b(g-a)=8.19\text{ N}."